A student weighs out a 6.12g sample of iron(II)iodide, transfers it to a 125mL volumetric flask, adds enough water to dissolve it and then adds water to the 125mL tic mark. What is the molarity of FeI2 in the resulting solution.
1
Expert's answer
2019-11-12T12:01:32-0500

Molarity (Сm) is the concentration of a solution expressed as the number of moles of solute per liter of solution.

Cm = n/V, n= m/M

m = 6,12g

M(FeI2) = 310 g/mol

V = 0,125 l

Cm = m/(M*V) = 6,12/(310*0,125) = 0,158 mol/l